OpenStudy (anonymous):

Can anyone help with Question 2 or 4 on this worksheet. I figured out the other problems. I will post document. 2)The graph to the right shows the rate of change in the temperature of a liquid within an experiment over a period of several hours. The temperature of the liquid at the beginning of this experiment (at t = 0) is 46degrees F. Use all this information to answer the questions that follow. 4)The graph at left of function, f , is drawn in such a way that each tick mark represents one unit. Give your best estimate for each of the following quantities:

OpenStudy (amistre64):

integration gives us the area under a curve

OpenStudy (amistre64):

min or max is determine when f' = 0 or is undefined

OpenStudy (amistre64):

f'(x) = x-8 ; from 0 to 4 = 8 ; from 4 to 8 f(x) = x^2/2 -8x + C ; from 0 to 4 = 8x from 4 to 8

OpenStudy (amistre64):

forgot a +C on tha tlast bit :)

OpenStudy (amistre64):

given that f(0) = 46 that amounts to f(x) = x^2/2 -8x + 46 we might be able to assume that 4 to 8 starts where the other one leaves off; but as is there is no sure way to define it